FEL Upgrade Project Weekly Brief

April 22-26, 2002




The completed and characterized "optical klystron" wiggler assembly for the 10 kW IR Upgrade project was successfully and safely delivered and placed in the FEL vault on Thursday, April 25. This completes one of the Upgrade project milestones. The Upgrade wiggler was a joint design and fabrication effort between JLab and two of our industrial partners Northrop Grumman and Advanced Energy Systems. Congratulations to all that worked on the wiggler including the project managers, including JLab's Steve Benson and George Biallas and Alan Todd from NG/AES.


Extra time was spent in the preparation and checking of the project monthly financial report for March since it is the midpoint of the fiscal year and this report is the basis of next week's (May 1-3) semiannual project review. The March monthly report was completed on Thursday and distributed to the ONR, AFRL and DOE contract monitors.

Presentations were finalized, briefing books were prepared and hardware displays were organized for next week's (May 1-3) review.

On Monday, April 22, Jefferson Lab hosted a visit by Mr. Kevin Cook, who is on committee staff for the House Energy and Water Committee. Following a briefing by Christoph Leemann, Mr. Cook toured the CEBAF accelerator site, Experimental Halls, FEL Facility and the Test Lab.

G. Neil also served this week Injector/Linac/ and Installation Panels on the Linac Coherent Light Source Project Review for the DOE to assess readiness for CD-1 submission for DOE approval to release Project Engineering Design funds. The $230M, 1.5 angstrom laser project at SLAC is aiming toward first light in 2007. Neil also presented a description of our laser to Nano Investing & Partnering Conference I, a full day conference at the NASA Training Center, Mountain View, CA in cooperation with the NASA-Ames Commercial Technology Office, the NASA-Ames Institute for Science Technology (NAIST), the NASA-Ames Center for Nanotechnology, and the NASA-Ames Research Park to set the groundwork to establish a joint government/privately-funded effort to commercialize FEL technology.

Congratulations are due to our intrepid FEL designer/builder extraordinaire: George-the-flash-Neil, who competed in last week's Boston Marathon and crossed the finish line standing up in 3 hrs and 9 minutes. Not bad for an aging boomer who spends the first 80 hours of each week on FELs.

WBS 4 (Injector Signed off drawings for the BPM situated between the light box and the buncher. Mounted the gun onto its stand.


Continued to work with W&M on refining the implantation process for the support tube on the gun.. The first test tube run showed still a few arcs, due possibly to silicon dioxide particulates falling from the chamber walls after nucleation within the N2/Si plasma and hitting the test tube. Thorough cleaning of the chamber will be performed from now on before each run. A new run is scheduled for this Friday after the installation of new oiless pumps.

Gun HVPS - A review of the Vendor's drawings for the HVPS Tank was done and approved. Fixtures to build and transport the conditioning resistor were completed this week. The conductive contacts will be applied this Friday and taken for firing early next week.

WBS 6 (RF):

Quarter HVPS - Wiring continues on both units. Unit 3 was tested for the control for the klystron heater voltage and current. It will be modified to make the adjustment less sensitive. The new PLC's are on order.

Quarter Klystrons - Work was started again on upgrading the LCW piping.

WBS 8 (Instrumentation):

A quick trip was made to Brookhaven National Lab to review the magnet power supply controls they are purchasing for the Spallation Neutron Source (SNS). The group who had done the RHIC controls wrote a specification, based on lessons learned, and had Apogee Systems design and build the interface. This consists of a VME power supply control board which connects to six power supply interface chassis via a pair of fiber optic cables. This provides 16 bits of digital I/O, 1-16 bit analog output, and 4-16 bit analog inputs with complete galvanic isolation to prevent ground loops. This complete system is operating under both EPICs and LabView. We will take advantage of a few man-years worth of effort and use this to interface to our new power supplies. One of these will be the 20 Amp, + 75 volt Danfsiks, these will supply the 12 'super trims'. This supply is being developed for the SNS project. The decision for the remaining 12 ~100 Amp supplies is still waiting on magnet damping measurements, this should be done in a week or so. A meeting was held to discuss the upcoming control upgrades to the existing trim racks and box power supplies, this should be quite clear with the work being complete by mid June. During these discussions Mark Augustine suggested moving the box power supply for arc 2, to the end of zone 1 instead of keeping them all together in zone 5. This was an excellent suggestion and was adopted, this will save >300 feet of (damn big) power cable (535mcm), ease the space constraints arising in zone 5 and ease the load on the cable trays. Not least of all the students won't have to drag it the length of the building.

The cable pulling continues, 15000 feet on ion pump power supply cable arrive this week, it will be pulled in on Tuesday. The Beam Position Monitor (BPM) Switched Electrode Electronics (SEE) cables were relocated, the 4-channel cables are being measured and will be done soon. Fabrication of the first new 8 channel stepper motor driver chassis is nearly complete. This will go to Al Grippo for software debugging next week. Fabrication of the 4 channel timing cards has also begun. Thanks to the operations crews for giving us a hand this week.

The prototype VME interface module (F0117) is now running in a test crate in lab 5 (iocfelt2). The F0117 module will be used to control a number of upgrade electronics including the BLMs (F0151), portions of the MPS system (F0118) and the drive laser's pulse control system (F0080 & more). The BLM break-out card (F0150) documentation has been checked out well and will be signed off in the next day or so allowing us to have a prototype made.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. The eighth coil is in potting. All coils are now available for the IR Machine’s four magnets.

• Upper and lower core’s plates final machining has finally started. We will examine their work on a vendor visit next Tuesday.

Injector Dipoles (DU/DV)

• Master Machine finished heat treatment of the core pieces and will start final machining next week. Most other parts are finished.

Arc Dipoles (GY, GX, GQ)

• Process Equipment Co. started manufacturing the cores and parts of the GQ-GX dipoles. They visited Jlab on Wednesday to discuss the fine points of machining the pole tips and were introduced to the pole tip changes we want to make.

• Wang NMR continues working on the winding fixture for the coils.

• AES will do one more run using their magnetic model of the GX dipoles at high energy with a larger shim only on the IR portion of the large face.

• Wang NMR continued winding the first GY coil.

• Bosma Machine has all GY core pieces machined and is now painting them. They will be ready for inspection on May 9.


3 inch quad (QX)

• The test stand was repaired and investigation of the effect of the Bang-Bang hysteresis protocol was being investigated.

Trim Quad (QT)

• Milhous Control of Virginia delivered the first article magnet. Workmanship was very good.

Sextupole (SF)

• DULY Research continues to work on completing drawings of the magnet.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• We are checking the design of the ARC Chambers.

• The second listing of stand positions was given to the Survey and Alignment Group.

• Girder assembly started.

• The injector chamber region is nearing completion

• Additional personnel were assigned to the recirculation region in front of the cryomodule.

• The optical chicane region with its chamber was taken up again after being interrupted for more urgent work.

WBS 10 (Wiggler):

The optical klystron was moved to the vault and positioned near its final location. It now awaits being surveyed into place and hooked up to water and power.

Analysis of data taken just before installation of the vacuum chamber indicates that Gauss level vertical fields indicated in the measurements were due to noise pickup in the probe as expected. The field outside the wiggler was measured and the difference between the measured and actual values was subtracted from the data. When this was done a field of up to 1.5 G was found in the wiggler. The cause of this was found to be an offset in the noise inside and outside the wiggler. When this is taken into account the DC vertical field in the wiggler falls to zero to within the level of the measurement uncertainty. When the source of the noise (the motor which drives the probe) is shut off, the vertical field drops to zero in the wiggler.

WBS 11 (Optics):

We have completed the rest of the drawing package for the optical cavity internals and it was signed today, and released for bid. Requisitions were signed for the additional optical transport gate valves in the user lab region, and for UHV gate valves for the optical cavity. Preliminary drawings for the optical cavity vacuum vessels will be ready for a design review next week.

Modifications to the existing mirror cassettes and design of the new turning cassettes continues. Design of the chamber and optics for the THz light pickoff continue. We ordered long lead items such as transducers for the laser beam dumps, optics for diagnostics, and a newer version pyroelectric camera that has better remote control capability.

This week, we determined that the Bessel filter introduced too much phase shift to be useful. We spoke to designers of similar systems at Duke and the ALS (at LBNL) and learned that using notch filters was the preferable method. Pspice models confirm this, and they have been added to the circuit and initial tests (made 4/25) look very promising. Work continues on the drive laser pulse compressor. We are installing a spectrometer to measure the bandwidth, and thus to determine the pulsewidth. Our tests the linear actuators for the optical cavity mirrors confirmed they produce well over the rated thrust of 25 lbs, we have not tried to make them stall. We will now procure the rest of the actuators. Thanks to the I&C group for setting up the necessary controls. We successfully Ni sputter-coated the edges of two mirrors (one sapphire, one ZnSe). These will now be tested for improved wetting of In:Ga eutectic, and improved adhesion of In:Ag braze. Thanks go to T. Elliott of the SRF Infrastructure Group, for his invaluable assistance.

Other activities:

We did more preparatory work to prepare for the May 3rd review.

We met with visiting scientist Malcolm Howell from LBL to discuss the design of the optical cavity, and how to best accommodate X-ray experiments.